### v1

# Explicit Motion Handling and Interactive Prompting for Video Camouflaged Object Detection

Abstract
Camouflage poses challenges in distinguishing a static target, whereas any movement of the target can break this disguise. Existing video camouflaged object detection (VCOD) approaches take noisy motion estimation as input or model motion implicitly, restricting detection performance in complex dynamic scenes. In this paper, we propose a novel Explicit Motion handling and Interactive Prompting framework for VCOD, dubbed EMIP, which handles motion cues explicitly using a frozen pre-trained optical flow fundamental model. EMIP is characterized by a two-stream architecture for simultaneously conducting camouflaged segmentation and optical flow estimation. Interactions across the dual streams are realized in an interactive prompting way that is inspired by emerging visual prompt learning. Two learnable modules, i.e. the camouflaged feeder and motion collector, are designed to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively, and enhance outputs of the both streams. The prompt fed to the motion stream is learned by supervising optical flow in a self-supervised manner. Furthermore, we show that long-term historical information can also be incorporated as a prompt into EMIP and achieve more robust results with temporal consistency. Experimental results demonstrate that our EMIP achieves new state-of-the-art records on popular VCOD benchmarks. The code will be publicly available.
1 Introduction
Camouflaged object detection (COD) aims at detecting and segmenting those hidden objects that exhibit high intrinsic similarity to their backgrounds. The inherent complexity of distinguishing camouflaged objects from their surroundings poses unique challenges when compared to general object detection [?] and salient object detection [?]. Recently, it has attracted interest of many researchers and facilities broad applications, e.g., medical image segmentation [??] and industrial inspection [??]. While significant progress has been made in image-based [???] and video-based [???] COD tasks, there still remains substantial room for development due to intrinsic difficulty of the tasks.
Figure 1: Different strategies of motion handling in VCOD: (a) Directly feed optical flow maps [??]; (b) Learn implicit motion cues and subsequently utilize them for mask decoding [?]; (c) The proposed interactive prompting paradigm handles motion cues explicitly using a pre-trained optical flow model, and simultaneously conduct optical flow estimation and camouflaged object segmentation. The fire/snowflake symbols denote that most of the model parameters are learnable/frozen in the proposed scheme.
Image-based COD approaches identify camouflaged targets using a static image. Typically, these targets exhibit strong visual resemblances to their backgrounds in terms of texture, color, or edges, posing a challenge for detection using appearance or geometric cues alone. Consequently, motion cues have been investigated in previous works [??] for video camouflaged object detection (VCOD) task. Fig. 1 summarizes the strategies adopted in handling motion cues by these previous works. As presented in Fig. 1 (a), off-the-shelf motion estimators (e.g., [??]) are directly employed for generating offline optical flow, serving as motion cues for identifying camouflaged objects. Notably, offline motion estimation, particularly in camouflaged scenarios, poses significant challenges for common optical flow estimators, often leading to noisy and inaccurate output optical flow. Such erroneous input could misguide detectors and thereby hinder overall performance. To address this issue, SLT-Net [?] proposes a distinct approach by implicit motion modeling in an online fashion (Fig. 1 (b)). While the overall motion modeling part is learnable, learning reliable motion of camouflaged objects from limited VCOD data may be intractable compared to using extensive training data for training common optical flow estimators [??]. Besides, due to the implicit nature of SLT-Net, no explicit regularization or evaluation can be adhered to the motion part, further making the learned motion less effective and reliable. Also, we believe that beyond motion, appearance cues are an important factor conducive to detection as well, since objects could be motion-free or still.
In this paper, we propose a novel Explicit Motion handling and Interactive Prompting framework for VCOD (Fig. 1 (c)), dubbed EMIP, which handles motion cues explicitly by freezing the upstream optical flow model mostly. EMIP is characterized by a two-stream architecture for simultaneously addressing camouflaged segmentation and optical flow estimation. Inspired by emerging advances of the visual prompt learning [???], we design an interactive prompting scheme to achieve interactions across the dual streams, as shown in Fig. 1 (c). Two modules, namely the camouflage feeder and motion collector, are designed to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively, and enhance outputs of the dual streams. Due to the absence of authentic optical flow for current VCOD datasets, we learn the prompt fed to the motion stream by supervising optical flow in a self-supervised manner. We also propose a long-term variant of EMIP by formulating historical features into the prompt to mitigate short-term prediction errors and further improve the accuracy. Benefiting from the above elaborate designs, the proposed EMIP  effectively leverages noise-robust motion to detect and segment those video camouflaged objects. The main contributions are summarized below:
  • We propose a novel framework for VCOD, dubbed EMIP, which handles motion cues explicitly using a frozen pre-trained optical flow fundamental model. EMIP is formulated into a novel two-stream architecture for simultaneously conducting camouflaged segmentation and optical flow estimation.
  • Inspired by visual prompt learning, the interactions across the two streams are realized in an interactive prompting way, and we propose two modules, i.e., the camouflaged feeder and motion collector, to incorporate segmentation-to-motion and motion-to-segmentation prompts, respectively.
  • We also propose a long-term variant of EMIP by formulating historical features into the prompt to mitigate short-term prediction errors.
  • EMIP together with its long-term variant, achieve new state-of-the-art records and outperform previous models by notable margins ($\sim$17.0%/5.5average improvement on F-measure/S-measure over the previous best model SLT-Net).
Figure 2: Overall architecture of the proposed EMIP, which consists of two separate streams: explicit motion modeling stream (upper) and object segmentation stream (lower). We use GMFlow  [?] as the fundamental model to handle motion cues. With the camouflage feeder and motion collector, segmentation and motion prompts are injected into each task-specific stream to compensate essential information. The fire/snowflake symbols indicate that the model parameters in this part or block are designated as learnable/frozen.
2 Related Work
Image-based COD. Methods in this type aim to discern camouflaged objects from a single RGB image. Early COD methods [??] relied on hand-crafted features to find targets hidden in the background. However, with the advent of deep learning, COD methods have undergone substantial advancement in recent years. Inspired by natural predatory behavior, SINet-V2 [?] and PFNet  [?] employed a coarse-to-fine strategy. They first generated a preliminary location map for camouflaged objects and then refined it for segmentation. To enhance performance, several studies integrated auxiliary task into a joint learning framework. MGL [?] combined classification or boundary detection task with COD. ZoomNet [?] employed a zooming in and out strategy to the original inputs and processed appearance features at three different scales. Jia et al. [?] proposed the SegMaR framework, an iterative refinement approach designed to locate, magnify, and detect camouflaged objects. HitNet [?] elevated low-resolution representations by leveraging high-resolution features in an iterative feedback loop, effectively mitigating challenges such as edge blurring and detail degradation.
Video-based COD. For the VCOD task, motion cues are crucial to camouflaged object detection. Bidau et al. [?] introduced a method by approximating various motion models derived from dense optical flow. Lamdouar et al.[?] introduced a video registration and segmentation network to detect camouflaged objects, employing optical flow and a difference image as inputs. However, the utilization of inaccurate optical flow may result in accumulative errors in mask prediction. To address this challenge, Cheng et al. [?] proposed a two-stage model that implicitly models and leverages motion information. Different from the above approaches, we propose to handle motion cues in an explicit way as aforementioned, and a two-stream architecture is designed to conduct both optical flow estimation and segmentation.
Visual Prompt Learning. Recently, prompt learning has emerged as a new paradigm that has significantly enhanced the performance of natural language processing (NLP) tasks [?]. Besides, the prompting paradigm has been adopted in many computer vision tasks [??]. Work [?] modified transformer layers by introducing memory tokens, constituting a set of learnable embedding vectors. VPT [?] employed a similar strategy by applying learnable embedding vectors to transformer encoders, achieving noteworthy performance across various downstream recognition tasks. Based on this idea, ViPT [?] integrated modality-complementary visual prompts for task-oriented multi-modal tracking. The popular Segment Anything Model (SAM) [?] integrates various visual prompts like points, boxes, or masks to achieve tailored object segmentation and exhibits decent zero-shot generalization. Previous researches have mainly focused on specific tasks like classification, tracking, or segmentation. In this paper, inspired by [?], we treat intermediate features from one stream as a prompt for injecting complementary information to the other, and propose an interactive prompting scheme for the VCOD task.
3 Methodology
In this work, we propose an end-to-end trainable network EMIP to jointly optimize camouflaged object detection and optical flow estimation. We input an adjacent frame pair $(I_{t},I_{t-1})$ of a video to EMIP, and output a binary segmentation mask of the reference frame $I_{t}$, together with optical flow estimation $\boldsymbol{V}$. The overall architecture of our EMIP is illustrated in Fig. 2. Note that, for object segmentation stream, a set of features $\{f_{n}^{i}\in\mathbb{R}^{H/{2^{i+1}}\times W/{2^{i+1}}\times C_{i}},n\in\{t,t-1\},i=1,...,4\}$ with different scales are extracted from a vision transformer backbone (i.e., PVTv2-B5 [?]), the same as SLT-Net [?], for fair comparison. $W$, $H$, and $C$ represent the width, height, and channel number, respectively. Similar to [?], we adopt the top-three features ($f_{t}^{2},f_{t}^{3},f_{t}^{4}$) for appearance modeling, and discard the first-layer feature $f_{t}^{1}$.
3.1 Fundamental Model for Motion Modeling
To achieve more effective integration of motion information for the VCOD task, we select GMFlow [?] as our fundamental model, which is trained on $\sim$50k frame pairs for optical flow estimation. As shown in Fig. 2, GMFlow can be delineated into two integral components, i.e., a CNN encoder and a transformer decoder. It initially employs the CNN encoder to capture low-level features such as edges, colors, and textures. Subsequently, the transformer decoder, composed of a sequence of self- and cross-attention layers, predicts optical flow map $\boldsymbol{V}\in\mathbb{R}^{H\times W\times 2}$ and matching distribution $\boldsymbol{M}\in\mathbb{R}^{H\times W\times H\times W}$ simultaneously. $\boldsymbol{V}$ is derived from $\boldsymbol{M}$ through operations such as pixel-grid sampling. Here, we flatten the first two dimensions of $\boldsymbol{M}$ to obtain $\boldsymbol{G}\in\mathbb{R}^{H\times W\times HW}$, and employ $\boldsymbol{G}$ as the motion prompt for interacting with segmentation features, leveraging its detailed pixel-to-pixel matching information. Additionally, the flow map $\boldsymbol{V}$ is utilized to reconstruct frame $I_{t}$ from $I_{t-1}$, leading to the computation of a self-supervised loss. For more details about our fundamental motion estimation model, we refer readers to GMFlow [?].
3.2 Segmentation-to-Motion Prompt
A good prompt can fully exploit the potential of the fundamental model, the same for our EMIP. Therefore, the segmentation prompt has to determine where and how to embed it in the fundamental model, considering the nature of the optical flow estimation task.
Position of prompt. The selection of the prompt position is guided by two key observations. Firstly, the choice of the fundamental model plays a pivotal role. Each fundamental model is tailored with a specific architecture; in the case of GMFlow within our EMIP framework, its transformer blocks synergistically form a functional unit crafted for calculating the similarity of pixel features between two frames. Thus, preserving the integrity of this architecture becomes paramount for the coherent generation of optical flow. Secondly, the architectural arrangement of neural networks involves a stratification where lower layers are dedicated to extracting basic features such as edges, colors, and textures, each exhibiting distinctive characteristics across various modalities [?]. This stratification has been shown by [?] to be particularly advantageous for task like optical flow estimation, emphasizing the importance of low-level visual correspondences. Leveraging this insight, we position the segmentation prompt at the lower layers of both the segmentation and motion modeling streams, with the direction being from segmentation to motion. Further details are discussed in Section 4.3, where we delve into various prompt position choices and their impact on results.
Camouflage feeder. To better incorporate the segmentation prompt into the motion stream, we design camouflage feeder. As shown in Fig. 2, our camouflage feeder is based on cross attention with a residual connection, which takes the flow input feature $f^{flow}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of query, and the segmentation prompt feature $f^{2}_{t}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ as the source of key and value. The residual connection is to maintain more query-related cues, and $d$ represents the embedding dimension. This process can be written as:
${\hat{f}}^{flow}_{t}={\overline{f}}^{flow}_{t}+\mathtt{GDFN}({\overline{f}}^{flow}_{t})$ (1)
where ${\overline{f}}^{flow}_{t}$ is formulated as:
${\overline{f}}^{flow}_{t}=f^{flow}_{t}+\mathtt{CA}(\mathbf{Q},\mathbf{K},\mathbf{V}),$ (2)
where $\mathtt{CA}$ is defined as:
$\mathtt{CA}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathtt{Softmax}(\mathbf{Q}\mathbf{K}^{\intercal}/\sqrt{d})\mathbf{V},$ (3)
and $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{\frac{HW}{8^{2}}\times d}$ represent query, key, and value matrices, respectively. $\bf Q/\bf K$ and $\bf V$ are derived by applying layer normalization and 3$\times$3 convolution to the input feature maps $f^{flow}_{t}$ and $f^{2}_{t}$, respectively. Here, we employ a Gated-Dconv Feed-forward Network (GDFN) [?] to suppress the aggregation of noise prompt. The GDFN process is defined as:
$\mathtt{GDFN}(\mathbf{X})=\mathtt{C1}\left[\phi(\mathtt{C3}(\mathtt{C1}(\mathtt{LN}(\mathbf{X}))))\odot\mathtt{C3}(\mathtt{C1}(\mathtt{LN}(\mathbf{X})))\right],$ (4)
where $\mathtt{C1}$ and $\mathtt{C3}$ are 1$\times$1 convolution and 3$\times$3 depth-wise convolution, respectively. $\phi$ is GELU function [?] and $\odot$ denotes element-wise multiplication. Note that, $\mathtt{GDFN}$ comprises two separate paths, one of which is activated with GELU function. Subsequently, an element-wise product is applied between these two paths. Literature [?] has demonstrated that the gating mechanism in GDFN can control information flow and yield better performance compared with the conventional feed-forward network (FN) [?] for feature restoration. For the same purpose, to mitigate the impact of the integrated noise prompts, we employ GDFN to control noise information flow. Finally, the resulting segmentation-prompted motion features $\hat{f}^{flow}_{t}$ with both knowledges of motion and appearance can be well adapted to subsequent motion reconstruction.
Prompt learning. To better learn the motion information, inspired by [?], we employ a self-supervised loss to optimize the motion-to-segmentation prompt learning. Further elaboration is provided in Section 3.4.
3.3 Motion-to-Segmentation Prompt
Motion-to-segmentation prompt formally provides an auxiliary motion flow that is both temporally and spatially coordinated with the segmentation stream. After obtaining the segmentation-prompted motion, we need to collect this knowledge back to the segmentation stream. Thus, we introduce another module named Motion Collector to integrate motion prompts into the segmentation stream. Given its functional equivalence of motion collector with the camouflage feeder, we opt to maintain the structural configuration of the interaction block rather than introducing a redesign. Specifically, upon receiving motion feature $\boldsymbol{G}$ (flattening the first two dimensions of $\boldsymbol{M}$), we commence by applying a 3$\times$3 convolution operation to $\boldsymbol{G}$ to adjust the channel dimensions to align with $f^{2}_{t}$. Then we feed the motion-to-segmentation prompt $\hat{\boldsymbol{G}}=\mathtt{Conv}(\boldsymbol{G})$ and $f^{2}_{t}$ to the motion collector to obtain the prompted appearance feature $f_{t\leftarrow t-1}^{2}$.
Subsequently, following that most VCOD works [??] use neighbor connection decoder (NCD) [?] to predict the segmentation map, we use NCD to decode the appearance features. NCD with camouflage feeder can be regarded as the prediction head for the motion fundamental model, similar to the concepts of some previous prompt learning methods [??] that design an extra prediction head after freezing the backbone parameters. Specifically, the appearance feature $f_{t\leftarrow t-1}^{2}$ prompted by motion combined with $f^{3}_{t}$ and $f^{4}_{t}$ are fed to NCD to obtain the predicted map. The optimization of motion-to-segmentation prompt and prediction head is under the constraint of segmentation loss, and more details will be given in Sec. 3.4.
Discussion. The motivation for our prompt learning strategy, freezing the motion stream and fully tuning the segmentation stream, is from the observed performance degradation of the PVT [?] pre-trained on ImageNet when it is directly applied to the VCOD task. Similar challenges are also encountered with large models like SAM [?], which struggle in camouflage scenes [?]. Currently, there lacks a freezable foundational model that excels on the VCOD task. In contrast, we find that the optical flow model could achieve good generalization in most VCOD scenarios, when most of its parameters are frozen.
3.4 Supervision and Loss Function
To ensure optimization for each component of the model, we separately define loss functions for the motion and segmentation stream, and then use these two losses to optimize the entire model jointly. Considering the lack of ground truth (GT) optical flow in video camouflaged scenarios, we propose a self-supervised strategy to learn the optical flow estimation stream. For the sake of brevity and clarity, we refer to [?] and warp the frame $I_{t-1}$ according to the optical flow estimation $\boldsymbol{V}$ to obtain the reconstructed reference frame for $I_{t}$, which is denoted as $\hat{I}_{t}$, i.e., $\hat{I}_{t}=\mathtt{Warp}(\boldsymbol{V},I_{t-1})$. We supervise the optical flow by computing the distance between $I_{t}$ and $\hat{I}_{t}$. Thus, the flow loss $\mathcal{L}_{\text{flow}}$ can be expressed as:
$\mathcal{L}_{\text{flow}}=\mathtt{SSIM}(I_{t},\hat{I}_{t}),$ (5)
where $\mathtt{SSIM}$ is pixel-wise photometric loss [?], which is commonly used for reflecting image distortion from three aspects: brightness, contrast, and structure.
For the segmentation stream, we employ a hybrid loss function [?], the same as SLT-Net [?]. The hybrid loss $\mathcal{L}_{\text{seg}}$ can be defined as $\mathcal{L}_{\text{seg}}=\mathcal{L}_{\text{IoU}}+\mathcal{L}_{\text{bce}}+\mathcal{L}_{\text{e-loss}}$, which includes IoU loss, binary cross-entropy loss, and enhanced-alignment loss. Finally, we jointly optimize the motion modeling and object segmentation streams, whose total loss is formulated as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{seg}}+\mathcal{L}_{\text{flow}}.$ (6)
3.5 Long-term Consistency Modeling
Figure 3: Long-term modeling scheme (${}^{\dagger}$).
The short-term model captures correlated information between adjacent frames, but due to the impact of noises and the limitation of motion estimation model, the short-term model may be less robust. To address this, long-term consideration can be further introduced to improve results’ consistency and therefore detection accuracy. To utilize long-term historical information, as well as maintain feasible computational cost, a dynamic memory model is designed to extract a long-term prompt, which formulates historical features into the prompt to mitigate short-term prediction errors.
Fig. 3 illustrates our long-term consistency modeling scheme, termed ${}^{\dagger}$. Specifically, during training, we freeze the short-term model after it converges and then add the long-term memory module for further learning. Given a video sequence {${{I}_{1},...,{I}_{t-1},{I}_{t}},t>1$}, we sequentially input current frame ${I}_{t}$ and preceding frame ${I}_{t-1}$ into the short-term model to obtain the appearance features $f^{2}_{t}$ and the motion-to-segmentation prompt $\hat{\boldsymbol{G}}$ between the two frames. The element-wise sum of $f^{2}_{t}$ and $\hat{\boldsymbol{G}}$ is conducted, and the results are then fed to an encoder consisting of two convolutional layers. The processed information is later stored into the memory pool. Considering the computational burden and minor impact of distant frames, we implement a FIFO (First-In First-Out) memory pool with fixed capacity ${L}$. Note that such a fixed capacity is configurable according to practical applications. Referring to the long-term setting of SLT-Net, we set ${L}$ to 5 in this paper. Since the memory pool stores appearance and motion features for both the current and historical frames, the features extracted from the current frame are then employed to query the memory pool using a space-time memory read block (STM) [?], yielding the long-term prompt ${P}_{L}$. Next, ${P}_{L}$ and $f^{2}_{t}$ are fed to the motion collector (re-use the same motion collector of the short-term period) for interaction, and the output of collector is subsequently fed to NCD for decoding.
We still adopt the hybrid loss $\mathcal{L}_{\text{seg}}$, the same as the short-term version, for training the long-term counterpart. It is worth noting that compared with SLT-Net which defines long-term modeling as a sequence-to-sequence problem (requiring all frames of a clip), our scheme utilizes only historical frames regarding the current frame. Hence, our scheme is theoretically more suitable for real-time and practical applications, where only past information is available.
4 Experiments
4.1 Implementation Details and Datasets
For fair comparisons, following the training strategy with [?], we train our model with a two-stage pipeline: first train the backbone on the static training set of COD10K (3,040 images) [?], and then fine-tune the whole model with the temporal components on the training set of MoCA-Mask (19,313 frames) [?]. Evaluation is performed on two widely recognized VCOD benchmarks: CAD (836 frames) [?] and MoCA-Mask test set (3,626 frames). Five evaluation metrics are adopted to assess model performance, i.e., structure measure ($\mathcal{S}_{\alpha}$[?], weighted F-measure ($F_{\beta}^{w}$[?], mean absolute error ($\mathcal{M}$), and mean value of Dice and IoU. The proposed EMIP is implemented by PyTorch [?] on a single NVIDIA 4090 GPU. During the training stage, the batch size is set to 6, and the maximum training epoch is 60. The entire model is optimized using the Adam optimizer [?] with a cosine annealing strategy. The maximum and minimum learning rates, along with the maximum adjusted iterations, are set to 1e-5, 1e-6, and 20, respectively. All the input images are resized to 352$\times$352 and subjected to data augmentation techniques, including color enhancement and random rotation. For inference, EMIP takes any two consecutive frames as input and outputs the segmentation prediction together with the corresponding optical flow estimation. However, for ${}^{\dagger}$, it requires consecutive-frame pairs to be input sequentially, in order to leverage its long-term property.
[…]
MoCA-Mask CAD
[…]
Image-based SINet [?] CVPR’20 0.598 0.231 0.028 0.276 0.202 0.636 0.346 0.041 0.381 0.283
SINet-V2 [?] TPAMI’22 0.588 0.204 0.031 0.245 0.180 0.653 0.382 0.039 0.413 0.318
DGNet [?] MIR’23 0.581 0.184 0.024 0.222 0.156 0.686 0.416 0.037 0.456 0.340
FEDER [?] CVPR’23 0.560 0.165 0.031 0.194 0.137 0.691 0.444 0.029 0.474 0.375
HitNet [?] AAAI’23 0.623 0.299 0.019 0.318 0.254 0.685 0.463 0.031 0.478 0.373
Video-based RCRNet [?] ICCV’19 0.555 0.138 0.033 0.171 0.116 0.627 0.287 0.048 0.309 0.229
MG [?] ICCV’21 0.530 0.168 0.067 0.181 0.127 0.594 0.336 0.059 0.368 0.268
PNS-Net [?] MICCAI’21 0.544 0.097 0.033 0.121 0.101 0.655 0.325 0.048 0.384 0.290
SLT-Net [?] CVPR’22 0.637 0.304 0.027 0.356 0.271 0.696 0.471 0.031 0.484 0.392
${}^{\dagger}$ [?] CVPR’22 0.631 0.311 0.027 0.360 0.272 0.696 0.481 0.030 0.493 0.401
EMIP (Ours) 0.669 0.374 0.017 0.424 0.326 0.710 0.504 0.029 0.528 0.415
${}^{\dagger}$ (Ours) 0.675 0.381 0.015 0.426 0.333 0.719 0.514 0.028 0.536 0.425
Figure 4: Visual comparisons of our models (EMIP and ${}^{\dagger}$) with eight state-of-the-art methods.
4.2 Comparisons with State-of-the-arts
To demonstrate the effectiveness of our method, we compare our model with 10 cutting-edge methods, comprising five image-based and five video-based. To ensure fair evaluation, predictions from these methods are either directly downloaded from the official website or generated using models retrained with open-source code.
Quantitative results are presented in Table 1. Notable performance gains can be observed of EMIP and ${}^{\dagger}$ over recently proposed image-based COD and VCOD models, like HitNet [?] and SLT-Net [?], across all five evaluation metrics. Several visual samples are shown in Fig. 4. EMIP appears to be more effective at simultaneously conducting motion and segmentation, and incorporating both enhanced features further makes it better for detecting camouflaged objects. Additionally, predictions of consecutive frames are listed in Fig. 6. One can see that incorporating the long-term modeling scheme, namely ${}^{\dagger}$, can further refine results and lead to boosted performance.
Table 2: Ablation results on camouflage feeder (CF) and motion collector (MC) modules of EMIP on MoCA-Mask dataset. The best results are highlighted in bold.
[…]
1 - - 0.627 0.268 0.031 0.340 0.254
2 - 0.657 0.349 0.020 0.394 0.300
3 0.669 0.374 0.017 0.424 0.326
4.3 Ablation Studies
We conduct thorough ablation studies by removing or substituting components from the complete EMIP.
Camouflage feeder and motion collector. We first validate the role of the two prompt integration modules, i.e., camouflage feeder and motion collector. As shown in Table 2, we start with the baseline model (#1), which directly uses the segmentation stream to extract single-frame information for prediction, i.e., removing the motion modeling stream of EMIP. Configuration #2 means using only motion information to prompt the segmentation stream, meanwhile discarding segmentation-to-motion prompt. Model #3 is the full EMIP. The integration of motion information is beneficial for the detection of camouflaged objects in video sequences. By incorporating the segmentation-to-motion prompt through the camouflage feeder into the motion stream, we enhance the robustness of motion cues, further resulting in improved performance for segmentation.
Full-tuning v.s. Freezing. Existing models usually address motion modeling by fully tuning all the parameters. In contrast, our EMIP adopts a prompt learning strategy to conduct motion modeling by freezing the motion fundamental model. From the comparisons in Table 3, where both settings are preloaded with pre-trained weights, the superiority of freezing the motion model over full-tuning is evident. This implies the prompt learning strategy for the motion modeling stream can better exploit its potential on limited VCOD data.
Table 3: Different learning strategies for the motion stream.
[…]
Full-tuning 0.645 0.326 0.018 0.364 0.281
Freezing (Ours) 0.669 0.374 0.017 0.424 0.326
Table 4: Different prompt destinations on the motion stream.
[…]
CNN Layer1 0.644 0.331 0.018 0.368 0.285
CNN Layer2 0.650 0.334 0.021 0.379 0.289
CNN Layer3 0.669 0.374 0.017 0.424 0.326
Prompt destination on the motion stream. The prompt destination should address where to feed the appearance prompt on the motion fundamental model. In this experiment, we utilize $f^{2}_{t}$ as both segmentation-to-motion prompt and motion collector input. Then we use it to prompt different CNN layers, and the evaluation results are presented in Table 4. It can be seen that prompting with the features after CNN Layer3 achieves superior performance. The underlying reason is that distinct-level features need the corresponding matching prompt to enhance level-specific representations.
Prompt source from the segmentation stream. To explore the most suitable appearance prompt to feed to the motion stream, we conduct another experiment with different prompt features, i.e., $f^{2}_{t}$, $f^{3}_{t}$, and $f^{4}_{t}$. The experimental results are presented in Table 5, where it is evident that utilizing appearance features $f^{2}_{t}$ as prompt yields the best performance. One possible explanation for this phenomenon is that slight motion of a moving object is likely to be missed in lower-resolution feature maps, namely $f^{3}_{t}$ and $f^{4}_{t}$, and therefore degrading the efficacy for prompting to the motion stream.
Table 5: Different prompt sources from the segmentation stream.
[…]
$f^{2}_{t}$ 0.669 0.374 0.017 0.424 0.326
$f^{3}_{t}$ 0.654 0.350 0.018 0.389 0.300
$f^{4}_{t}$ 0.624 0.289 0.023 0.331 0.251
Prompt destination on the segmentation stream. We leverage motion information to selectively prompt distinct appearance features or simultaneously prompt all features by introducing additional motion collectors. The results, as presented in Table 6, demonstrate that utilizing motion to prompt feature $f^{2}_{t}$ outperforms other features ($f^{3}_{t}$ or $f^{4}_{t}$) or all features collectively. This implies that semantic features at corresponding hierarchical levels between segmentation and motion tasks can mutually enhance the representation of the original features. Yet, the integration of features from disparate levels, which lack alignment, may compromise the fidelity of the original feature representation.
Table 6: Different prompt destinations on the segmentation stream.
[…]
0.669 0.374 0.017 0.424 0.326
0.642 0.319 0.019 0.359 0.271
0.646 0.325 0.021 0.366 0.283
0.639 0.313 0.020 0.361 0.271
Effectiveness of motion self-supervision. Quantitative and qualitative results are reported to validate the effectiveness of different training strategies for the motion stream. Due to the absence of optical flow ground truth, conducting quantitative analyses for the output optical flow is not feasible. Thus, we evaluate its effectiveness via segmentation/prediction performance. From Table 7, the model with a self-supervised loss achieves better performance, which demonstrates that the prompted motion under self-supervision is beneficial to boosting detection performance.
Fig. 5 visualizes comparisons of our model with different settings in terms of optical flow prediction for camouflaged scenarios. The results show that sometimes the original GMFlow cannot perceive camouflaged moving objects (Fig. 5 (c)), whereas ours that incorporates camouflage information with the prompt learning paradigm can better detect the targets (Fig. 5 (a)). Fig. 5 (d) further shows that the lack of self-supervised loss in our EMIP leads to worse flow prediction. In contrast, full-tuning the GMFlow part (Fig. 5 (b), correspond those results in Table 3) is hardly generalized in new camouflaged scenarios.
Figure 5: Visual comparisons employing different designs: (a) Our self-supervision (default EMIP), (b) Full-tuning, (c) GMFlow, (d) Ours w/o self-supervised loss.
Table 7: EMIP without/with the self-supervised loss.
[…]
w/o self-supervised 0.644 0.321 0.020 0.363 0.277
w/ self-supervised 0.669 0.374 0.017 0.424 0.326
Effectiveness of long-term modeling. As aforementioned in Table 1, by introducing the long-term consistency prompt to the short-term model EMIP, we have further enhanced the performance on all five metrics. Additionally, visual comparisons on the same clip in Fig. 6 show that the incorporation of long-term consistency effectively harnesses temporal consistency, thereby reducing errors in short-term prediction.
Figure 6: Visual comparisons on consecutive video frames.
5 Conclusion
We propose EMIP, an innovative framework for VCOD that explicitly handles motion cues through the utilization of a frozen pre-trained optical flow fundamental model. EMIP adopts a novel two-stream architecture, concurrently addressing camouflaged segmentation and optical flow estimation. The core idea of interaction between these two streams is orchestrated through an interactive prompting mechanism. Experimental results show the paradigm of interactive prompting of EMIP can enhance outputs of both streams, further achieving accurate prediction. In addition, we present an extended version of EMIP, incorporating historical features into the prompt to alleviate short-term prediction errors and enhance overall accuracy. Our contributions not only achieve compelling results on two VCOD benchmarks but also provide fresh insights into addressing the challenging VCOD task. We hope that the proposed framework will serve as a catalyst for inspiring further research in this emerging field.